# Quantifying the Uniqueness of Donald Trump in Presidential Discourse

### v1

@@ 3.3 References to political opponents in speech
Rates of opponent mentions
@@
Since the opponent mentions are more common in presidential debates, we revisit our LLM-based measure for the debates to confirm that Trump is not unique simply because he is more likely to mention opponents. Fig. 6 shows that the LLM-based uniqueness score rankings are relatively consistent across opponent and non-opponent mentions. Moreover, Trump’s language is consistently unique in these debates, whether or not he calls out an opponent.
Figure 6: Uniqueness of speech broken down by opponent mentions, for debates (the error bars represent 95%-confidence intervals). Trump is most distinctive regardless of whether he references an opponent or not.
Fightin’ Words adjective overlap
To identify which adjectives are most strongly associated with describing opponents for each candidate, we calculate the odds ratios between opponent mention sentences and non-opponent mention sentences, i.e., Fightin’ Words (FW) [21]. After these odds ratios are calculated, we compare the words most commonly associated with opponent mentions across candidates and calculate an overlap score. This FW overlap metric for each candidate is calculated as the average number of speakers who share each of the candidate’s top-$n$ FW adjectives. See the Materials and Methods section for further details on the construction of this metric.
@@
(a) FW overlap scores across different top-$n$ thresholds, comparing Trump and the other candidates aggregated by party.
(b) Top-15 FW overlap score in debates for each candidate
Figure 7: Fightin’ Words overlap results. Trump’s FW associated with opponent mentions generally have the lowest overlap in adjective usage compared to other candidates, which is another indication that his language is distinctive.
Mentions Opponents: Y
Mentions Opponents: N
left (2.07), long (1.85), tough (1.80), own (1.79), ok (1.79), bigger (1.79), worse (1.66), radical (1.57), super (1.50), real (1.50), effective (1.47), xenophobic (1.46), liberal (1.46), disgraceful (1.46), massive (1.33), single (1.33), political (1.09), last (1.09), economic (1.08), short (1.08), huge (1.04), various (1.04), red (1.04), upset (1.04), back (1.04)
good (-3.63), great (-3.25), important (-2.36), more (-2.08), inner (-2.00), strong (-1.99), right (-1.84), proud (-1.80), other (-1.79), expensive (-1.64), old (-1.53), big (-1.53), better (-1.50), greatest (-1.44), young (-1.41), sad (-1.41), able (-1.31), much (-1.31), african (-1.27), fine (-1.27), beautiful (-1.27), tougher (-1.18), nice (-1.18), least (-1.12), sure (-1.12)
@@
4 Qualitative analysis of unique political speech
Having found that Trump’s speech is measurably distinct from other politicians, we now further investigate the sources of these differences. Our LLM-based score enables us to identify novel patterns in Trump’s speech. From qualitative analysis of the most and least unique (as determined by our LLM-based score) sentences from Trump, we find that Trump’s speech is distinctive not merely for the particular words that he uses, but for his patterns of repetition and short sentences.
Unique Patterns of Repetition
Here, we focus on a subset of Trump’s speech that reveals a distinctive pattern — his proclivity to repeat a single word for emphasis. In the following examples, a ${}^{\dagger}$ denotes the sentence whose uniqueness score is under assessment. The preceding sentences form the context that the LLM evaluates when scoring the sentence.
The following exchange from a debate between Donald Trump and Hillary Clinton results in one of Trump’s most unique utterances (score 0.23):
Clinton: And when we talk about your business, you’ve taken business bankruptcy six times.
Clinton: There are a lot of great business people that have never taken bankruptcy once.
Clinton: You call yourself the King of Debt.
Clinton: You talk about leverage.
Clinton: You even at one time suggested that you would try to negotiate down the national debt of the United States.
Trump: Wrong.
Trump: ${}^{\dagger}$
This emphatic repetition of “wrong” turns out to be unique to Trump: he is the only candidate to interject “wrong”, which he does 13 times overall in the debates. Furthermore, he specifically repeats the word twice in succession on six separate occasions in his debates, an utterance that is not shared with any other candidate.
This habit of repeating a single word is not confined to the debate setting. One of Trump’s most unique sentences from SOTU speeches include the following from 2019:
Trump: My budget will ask Democrats and Republicans to make the needed commitment to eliminate the HIV epidemic in the United States within 10 years.
Trump: We have made incredible strides.
Trump: ${}^{\dagger}$
Additionally, in 2020:
Trump: I’ve also made an ironclad pledge to American families: We will always protect patients with preexisting conditions.
Trump: And we will always protect your Medicare, and we will always protect your Social Security.
Trump: ${}^{\dagger}$
These examples suggest that a distinguishing feature of Trump’s speech is the use of short phrases that are repeated for emphasis. In fact, Trump has the highest raw count of one-word sentences in all three data types and the highest frequency of one-word sentences in the SOTU and campaign datasets.
@@ 5 Discussion
Whether on the campaign trail, the debate stage, or the official dais of the House of Representatives, Trump speaks differently from all modern presidential candidates. We confirm his distinctiveness for speaking in shorter, simpler, and more repetitive sentences. We further quantify the uniqueness of his overall speech compared to that of other presidential candidates. Finally, we demonstrate that Trump uses language that is more divisive, antagonistic, and explicitly focused on his political opponents.
@@
The differences between Trump and other candidates do not appear to be an artifact of secular communication trends (see §B.3), whether in the general coarsening of public discourse or the tendency to speak in simpler sentences. Across multiple points of comparison, observed differences between Trump and other contemporary candidates appear every bit as large as those between Trump and candidates from the 1960s and 70s.
@@
Our findings, of course, come with a variety of limitations. The campaign data, for instance, include samples from only the most recent candidates and exclude certain communication formats, such as interviews. None of our datasets cover instances when surrogates (family members, vice presidents, etc.) speak on behalf of a candidate. While other works have examined Trump’s tweets [22], we do not assess social media postings, due to the lack of comparative data for other candidates, especially older ones. Furthermore, our data collection ends at 2022, excluding more recent remarks from Trump that have been flagged in the media for their escalated divisiveness [23]. While investigating new dimensions of political speech, we recognize that others — e.g., its analytical sophistication or propensity to endorse racial, gender, or sexual stereotypes — warrant still further scrutiny. Also, as previously noted, all of our metrics come with tradeoffs.
@@
Still, the tools we develop can be applied to a wide variety of settings. While we use them to compare speech patterns among presidential candidates, they can just as easily be deployed to analyze the language used by any public officials. And in addition to characterizing overall differences in speech patterns, we illuminate ways of assessing particular qualities that appear in either stand-alone speeches or interactive exchanges.
@@
The substantive findings presented herein, moreover, establish the overwhelming uniqueness of Trump’s speech patterns, just as they reveal particular qualities that distinguish Trump from all modern US presidential candidates. More research is required to map these speech patterns into larger political strategy. Nonetheless, we conjecture that these qualities broadly contribute to Trump’s enduring appeal as a populist who unabashedly denounces established political enemies in a historical period of acute polarization, distrust, and division.
@@ A.1 Data collection
    All campaign documents from 1932-2020 are collected initially. We manually identify keywords that indicate campaign speeches of interest and filter speech titles based on those keywords. Keywords selected are: “remarks”, “speech”, and “address”. Manually removed documents include: speeches with reporters, question/answer format speeches, press releases, town halls, press releases, co-appearances with other politicians and spouses, etc.
@@
    Following these filtering steps, our final campaign dataset includes only speeches delivered within the month before election day and only from candidates since 2008. This smaller subset is used to generate the results presented in the main paper.
@@ A.2 Sentence and opponent mention tagging
    After each dataset is filtered, sentences are tagged as mentioning an opponent or not through an automated process. Sentences are classified as definitely including an opponent mention, possibly including an opponent mention, or not including an opponent mention based on keywords and the presence of parts of speech. For debates, “opponents” are identified by the name(s) of the debate partner or their party. After automatically tagging the debates dataset, the sentences that are labeled as possibly including an opponent mention are manually reviewed by an expert team of four researchers. The team compares a subset of overlapping ratings for consistency. The Cohen’s kappa for inter-coder agreement is roughly 0.8, which indicates substantial agreement.
    For SOTU, opponent mentions only include those of other presidential candidates; names of candidates who never held presidential office are not categorized as mentions of opponents, in contrast to the debates dataset. Following a manual review of selected speeches from the SOTU dataset, it is concluded that references to non-presidential figures are much more neutral and considerably less frequent in State of the Union speeches than they are in debate contexts.
@@ A.3 Language model training and analysis
    We fine-tune a pre-trained GPT-2 model using the huggingface [24] and PyTorch Lightning [25] libraries for each of the three data types, resulting in three different model types:
@@
    To preprocess data for training, we parsed each speech into sentences, prefixed each sentence with the speaker prompt (e.g., “Donald Trump:”, and masked any named entities (as identified by spaCy NER tagger) with a <ENT> mask. [^6] Versions of each model were also trained on unmasked data. The results from these models are consistent with our main findings and are presented in §B.5. We fine-tune each model for 10 epochs on all available corresponding data with a learning rate of $5e$-$5$.
@@ A.5 Divisiveness lexicon
    Ten seed “divisive” terms are chosen manually by NLP and political science experts as common politically divisive English terms, before seeing the actual speech data. The ten seed terms used are: “stupid”, “dishonest”, “unamerican”, “idiot”, “deplorable”, “pathetic”, “immoral”, “disgrace”, “incompetent”, “foolish”. With these seed terms, 350 additional terms with the highest cosine similarity in the vector space model are added to the lexicon. This initial collection of 360 terms is analyzed by four researchers who annotate each term as “divisive”or “not divisive” based on the following criteria:
@@
    • Should not include words that can often be used both divisively and non-divisively
@@
    Only words that receive a majority of votes by the annotators are included in the final lexicon. The average pairwise Cohen’s Kappa agreement score of annotations is 0.541, which indicated a moderate level of inter-coder agreement. The final size of the lexicon is 178 words. The entire lexicon can be found in §C.
@@ A.6 Fightin’ Words overlap metric
    Monroe et al. [21] introduce a methodology for lexical feature selection that calculates odds ratios between of word probabilities between two related corpora. We use this method with the informative Dirichlet prior to obtain the Fightin’ Words in a data type for each candidate.
    For each candidate $s\in C$, we specifically get the set of Fightin’ Words between the words associated with their opponent mentions (${FW}_{Y}(s)$) versus the set of those that are not (${FW}_{N}(s)$). We then examine the top-$n$ of these sets respectively, ${FW}_{Y}^{n}(s)$ and ${FW}_{N}^{n}(s)$. For $Y$ opponent mentions, we create a graph representation as follows (see Fig. 8):
@@
    • word nodes $W$, where $W$ corresponds to the union of each candidate’s top-$n$ (${FW}_{Y}^{n}(s)$), i.e., $W=\bigcup_{s\in C}{FW}_{Y}^{n}(s)$
    • edges $E(S,W)$, where edge $e$ is added between $s\in S$ and $w\in W$ if $w$ is in $FW_{Y}^{n}(s)$
@@
    Equivalently, the overlap metric for $s$ can be thought of as the average number of speaker sets $FW_{Y}^{n}(c_{i}),~{}\forall c_{i}\in C$ that the top-$n$ words of $s$ appear in. A low $OM_{Y}^{n}(s)$ indicates that speaker $s$ uses more distinct language to refer to opponents, while a higher score corresponds to opponent-referring language that is similar to that of other candidates.
    Appendix B Additional LM-based Uniqueness Results

### v2

@@ 3.3 References to political opponents in speech
Rates of opponent mentions.
@@
Since the opponent mentions are more common in presidential debates, we revisit our LLM-based measure for the debates to confirm that Trump is not unique simply because he is more likely to mention opponents. Fig. 6 shows that the LLM-based uniqueness score rankings are relatively consistent across opponent and nonopponent mentions. Moreover, Trump’s language is consistently unique in these debates, whether or not he calls out an opponent. Concurrently, sentences that refer to an opponent tend to be more distinct than those that do not (see §E). And in debates, we additionally find that sentences that mention opponents contain significantly higher frequencies of divisive words. If you are looking to isolate the distinguishing characteristics of a candidate’s speech patterns, therefore, you would do well to focus on how they talk about their political opponents.
Figure 6: Uniqueness of speech broken down by opponent mentions, for debates (the error bars represent 95% CI). Trump is most distinctive regardless of whether he references an opponent or not.
Fightin’ Words adjective overlap.
To identify which adjectives are most strongly associated with describing opponents for each candidate, we calculate the odds ratios between opponent mention sentences and nonopponent mention sentences, i.e., Fightin’ Words (FW) [21]. After these odds ratios are calculated, we compare the words most commonly associated with opponent mentions across candidates and calculate an overlap score. This FW overlap metric for each candidate is calculated as the average number of speakers who share each of the candidate’s top-$n$ FW adjectives. See §A.6 for further details on the construction of this metric.
@@
Figure 7: FW overlap scores across different top-$n$ thresholds, comparing Trump and the other candidates aggregated by party.
Figure 8: Top-15 FW overlap score in debates for each candidate. Trump’s FW associated with opponent mentions generally have the lowest overlap in adjective usage compared to other candidates, which is another indication that his language is distinctive.
@@
Mentions Opponents: Y Mentions Opponents: N
left (2.07), long (1.85), tough (1.80), own (1.79), ok (1.79), bigger (1.79), worse (1.66), radical (1.57), super (1.50), real (1.50), effective (1.47), xenophobic (1.46), liberal (1.46), disgraceful (1.46), massive (1.33), single (1.33), political (1.09), last (1.09), economic (1.08), short (1.08), huge (1.04), various (1.04), red (1.04), upset (1.04), back (1.04) good (-3.63), great (-3.25), important (-2.36), more (-2.08), inner (-2.00), strong (-1.99), right (-1.84), proud (-1.80), other (-1.79), expensive (-1.64), old (-1.53), big (-1.53), better (-1.50), greatest (-1.44), young (-1.41), sad (-1.41), able (-1.31), much (-1.31), african (-1.27), fine (-1.27), beautiful (-1.27), tougher (-1.18), nice (-1.18), least (-1.12), sure (-1.12)
@@ 4 Discussion
In this work, we proposed novel approaches to quantify the uniqueness and divisiveness of presidential discourse. Our results show that whether on the campaign trail, the debate stage, or the official dais of the House of Representatives, Trump speaks differently from all modern presidents and presidential candidates. We confirm his distinctiveness for speaking in shorter, simpler, and more repetitive sentences. We further quantify the uniqueness of his overall speech compared to that of other presidential candidates. Finally, we demonstrate that Trump uses language that is more divisive, antagonistic, and explicitly focused on his political opponents.
@@
The differences between Trump and other candidates do not appear to be an artifact of secular communication trends (see §B.3), whether in the general coarsening of public discourse or the tendency to speak in simpler sentences. Across multiple points of comparison, observed differences between Trump and other contemporary candidates appear every bit as large as those between Trump and candidates from the 1960s and 1970s.
@@
Our findings, of course, come with a variety of limitations. The campaign data, for instance, include samples from only the most recent candidates and exclude certain communication formats, such as interviews. None of our datasets cover instances when surrogates (family members, vice presidents, etc.) speak on behalf of a candidate. While other works have examined Trump’s tweets [25], we do not assess any of his social media postings for lack of comparable data for other candidates, especially older ones. Furthermore, our data collection ends at 2022, excluding more recent remarks from Trump that have been flagged in the media for their escalated divisiveness [26].
Our research invites a variety of extensions. Given that values of the LLM-based uniqueness metric can only be understood relative to the elements of a selected sample, future work would do well to compare Trump to other speakers, such as populist leaders in other countries. Just as we investigate divisiveness as a new dimension of political speech with particular resonance for Trump’s rise to power, other scholars might evaluate other dimensions of political speech, such as its analytical sophistication or propensity to endorse racial, gender, or sexual stereotypes. Additionally, future research might explore the possibility of contagion effects from Trump’s speech patterns to other political figures, as politicians periodically emulate and influence one another.
@@
We also recognize that all of our metrics come with tradeoffs, which we have elucidated throughout this essay. Still, the tools we develop can be applied to a wide variety of settings. While we use them to compare speech patterns among presidents and presidential candidates, they can just as easily be deployed to analyze the language used by any public officials. And in addition to characterizing overall differences in speech patterns, we illuminate ways of assessing particular qualities that appear in either stand-alone speeches or interactive exchanges.
@@
The substantive findings presented herein, moreover, establish the uniqueness of Trump’s speech patterns, just as they reveal particular qualities that distinguish Trump from all modern US presidential candidates. More research, of course, is required to map these speech patterns into larger political strategy. Nonetheless, we conjecture that these qualities broadly contribute to Trump’s enduring appeal as a populist who unabashedly denounces established political enemies in a historical period of acute polarization, distrust, and division.
Acknowledgments
This work is in part supported by a Data & Democracy seed funding at the University of Chicago.
Data availability
The data underlying this article are available in the American Presidency Project at http://www.presidency.ucsb.edu/ws, and can be accessed publicly. For reproducibility, we release our data and code at https://github.com/ChicagoHAI/quantifying-unique-and-divisive-speech.
Appendix A Implementation Details
@@ A.1 Data collection
All campaign documents from 1932 to 2020 are collected initially. We manually identify keywords that indicate campaign speeches of interest and filter speech titles based on those keywords. Keywords selected are: “remarks”, “speech”, and “address”. Manually removed documents include: speeches with reporters, question/answer format speeches, press releases, town halls, press releases, co-appearances with other politicians and spouses, etc.
@@
After filtering, our final campaign dataset includes only speeches delivered within the month before election day and only from candidates since 2008. This smaller subset is used to generate the results presented in the main paper.
@@ A.2 Sentence and opponent mention tagging
After each dataset is filtered, sentences are tagged as mentioning an opponent or not through an automated process. Sentences are classified as definitely including an opponent mention, possibly including an opponent mention, or not including an opponent mention based on keywords and the presence of parts of speech. For debates, “opponents” are identified by the name(s) of the debate partner or their party. After automatically tagging the debates dataset, the sentences that are labeled as possibly including an opponent mention are manually reviewed by an expert team of four researchers. The team compares a subset of pairwise overlapping ratings for consistency. The Cohen’s $\kappa$ for inter-coder agreement is roughly 0.8, which indicates substantial agreement.
For SOTU, opponent mentions only include those of other presidential candidates; names of candidates who never held presidential office are not categorized as mentions of opponents, in contrast to the debates dataset. Following a manual review of selected speeches from the SOTU dataset, it is concluded that references to nonpresidential figures are much more neutral and considerably less frequent in State of the Union speeches than they are in debate contexts.
@@ A.3 Language model training and analysis
We fine-tune a pretrained, 124M parameter GPT-2 model using the huggingface [22] and PyTorch Lightning [27] libraries for each of the three data types, resulting in three different language models:
@@
To preprocess data for training, we parsed each speech into sentences, prefixed each sentence with the speaker prompt (e.g., “Donald Trump:”, and masked any named entities (as identified by spaCy NER tagger) with a <ENT> mask. [^7] Versions of each model were also trained on unmasked data. The results from these models are consistent with our main findings and are presented in §B.5.) We fine-tune each model for 10 epochs on all available corresponding data with a learning rate of $5e$-$5$. Validation with larger, more recent LLMs is shared in §B.6.
@@ A.5 Divisiveness lexicon
Ten seed “divisive” terms are chosen manually by NLP and political science experts as common politically divisive English terms, before seeing the actual speech data. The ten seed terms used are: “stupid”, “dishonest”, “unamerican”, “idiot”, “deplorable”, “pathetic”, “immoral”, “disgrace”, “incompetent”, “foolish”. With these seed terms, 350 additional terms with the highest cosine similarity in the vector space model are added to the lexicon. Each of these initial 360 terms is analyzed by four researchers as “divisive”or “not divisive” based on the following criteria:
@@
  • Should not include words that can often be used both divisively and nondivisively
@@
Only words that receive a majority of votes by the annotators are included in the final lexicon. The final size of the lexicon is 178 words. The entire lexicon can be found in the §C.1.
Annotator agreement.
Following [30], in order to analyze how often the annotators agree with each other, we additionally calculate the percentage of times the majority class has size 4 (all annotators agree), size 3 (all but one agree), and size 2 (even split). Table 5 shows these agreement values: 58.9% of the 360 initial terms had full annotator agreement, and 86.7% of these 360 have at least 3/4 annotators agreeing on the label. When limiting to the 178 terms that at least 3 out of 4 annotators agree are divisive, 69.1% have full agreement on the label (and 100% have majority agreement, by definition). So, the final terms include primarily instances that have high agreement scores.
Table 5: Percentage of majority class annotator agreement on binary labels (divisive? “yes” or “no”). 58.9% of the 360 initial terms had full annotator agreement, and 86.7% of these 360 have at least 3/4 annotators agreeing on the label. When limiting to the 178 terms that at least 3/4 annotators agree are divisive, 69.1% have full agreement on the label.
Majority class size (%)
Terms four three two
360 initial 58.9 27.8 13.3
178 final 69.1 30.9 0.0
Since there are four annotators, we calculate Fleiss’ $\kappa$ for interannotator agreement, as opposed to Cohen’s $\kappa$ which is designed for two raters. For the initial 360 terms, we obtain a Fleiss’ $\kappa$ of $0.54$, which indicates a moderate level of agreement. Discussion of this score and why it may underestimate true agreement is included in §C.5.
@@ A.6 Fightin’ Words overlap metric
Monroe et al. [21] introduce a methodology for lexical feature selection that calculates odds ratios between of word probabilities between two related corpora. We use this method with the informative Dirichlet prior to obtain the Fightin’ Words (FW) in a data type for each candidate.
[…]
@@
[…]
@@
[…]
Appendix B LLM-based Uniqueness Robustness Checks
B.1 BPC
Recall that we employ bits-per-character (BPC), also known as bits-per-byte, as a measure of “predictability” of a sentence. Lower BPC values correspond to higher predictability. We use BPC instead of perplexity or loss directly to account for variation in tokenization techniques. BPC is the sum of cross-entropy losses of each token in a sentence, divided by the number of characters (bytes) in the sentences. Fig. 10 shows average sentence BPC per candidate (Fig. 10(a)) and aggregated by party over sentence length (Fig. 10(b)). While in campaigns, Trump is the most unpredictable among those candidates, he is ranks closer to the middle for predictability in debates and SOTU.
B.2 Correlation with Readability Metrics
We calculate the Spearman correlation coefficients for our uniqueness metric with existing readability scores. Our metric has little to no correlation with readability, while different readability indexes show strong correlation with each other (Fig. 11). Fig. 11 also confirms that our metric is dintinct from sentence length.
B.3 Uniqueness scores over time
B.4 Top %-ile uniqueness
B.5 LLM-based uniqueness results with unmasked model
B.6 Validation with other LLMs
We validate our uniqueness metric implementation with other LLMs. Specifically, we recalculate scores with the Gemma-2B and Phi-1.5 (1.3B) base models [23, 24]. These models are considerably larger than the 124M parameters of GPT-2 [22]. Like our original setup up for GPT-2, we further train each model on all available corresponding data with a learning rate of $5e$-$5$. Each model is trained with LoRA tuning for 15 epochs, using LLaMA-Factory [31]. As shown in Fig. 18, while the exact scores differ, the trends are consistent with the main results: that is, Trump’s speech is the most unique among the selection of Democratic and Republican presidential candidates.
Appendix C Additional Divisive Word Lexicon Results
C.1 Divisive Word Lexicon
Table 6 contains the 178 words in our proposed divisive word lexicon.
C.2 Usage over time
C.3 Usage by candidate
C.4 Utterances Using Divisive Words
C.5 Annotator Agreement
For the initial 360 terms, we obtain a Fleiss’ $\kappa$ of $0.54$, which indicates a moderate level of agreement. Since divisiveness is a somewhat subjective phenomenon, this moderate agreement is similar to that obtained on other subjective lexicons, such as the word-emotion association (average Fleiss’ $\kappa$ of 0.29) [32] and the initial terms in the polarization dictionary (Cohen’s $\kappa$ of 0.61) [17]. Our final set of 178 terms has full agreement from the majority of annotators, but we still discuss the implications of our Fleiss’ $\kappa$ below.
Agreement is affected by several characteristics of divisiveness. Firstly , divisiveness is a spectrum and terms can be “a little” to “extremely” divisive. Annotators only give binary labels, so disagreement may be higher due to lack of nuance. Relatedly, the words are presented out of context; they may be more or less divisive depending on usage. However, identifying all possible contexts of our lexicon usage is costly and nontrivial. Finally, like the word-emotion association lexicon [32], our initial data is biased, with more “not divisive” terms than “divisive” terms (based on only 31.6% of the 360 terms receiving unanimous agreement). Both Cohen’s and Fleiss’ $\kappa$ are considered conservative in such settings with label imbalance [32, 33, 34]. As such, these $\kappa$ values may underestimate the true rater agreement for these cases.
Ultimately, we introduce the first such lexicon of divisiveness for text analysis. Usage of our lexicon may be augmented with additional strategies like dependency parsing and leveraging surrounding context.
Appendix D Additional Opponent Mention Results
D.1 Rate of opponent mentions
D.2 Fightin’ Words overlap
Appendix E Relating uniqueness, divisiveness, and opponent mentions.
Fig. 23(a) shows that in debates and campaign speeches, sentences that use divisive language tend to be more unique as well. The Spearman correlation coefficient between divisiveness and uniqueness is 0.01 for debates and campaigns ($p<$0.05 in campaigns and $p=$0.30 in debates). For sentences containing opponent mentions in debates, Fig. 23(b) shows that such utterances tend to be more distinctive and Fig. 23(c) shows that they tend to have higher divisive word usage as well. The Spearman correlation for opponent mentions is 0.05 and 0.02 ($p<$0.05) between uniqueness and divisiveness, respectively.